OpenStudy (anonymous):

I can't use a calculator for this question so i need HELP. find the x-intercepts of f(x)= x^3- 2x^2 -x +2

OpenStudy (agentnao):

Ah! This is easy peasy my dear :) Factor by grouping and then find the zeros!

OpenStudy (agentnao):

If you need me to be more in depth, I can show you step by step.

OpenStudy (anonymous):

yes please, step by step would be nice :)

OpenStudy (anonymous):

.....

OpenStudy (agentnao):

Sure! x^3 - 2x^2 - x + 2 = 0 (x^3 - x) (- 2x^2 + 2) = 0 <-- All I did here was rearrange the terms & put them in ( ) x(x^2 -1) -2(x^2 - 1) = 0 <-- Factor (x - 2)(x^2 -1) = 0 <-- Just grouping again x - 2 = 0 --> x = 2 <-- Setting them equal to zero x^2 - 1 = 0 --> x^2 = 1 <-- Now square root both sides x = 1 Your intercepts are x = 1 and 2 :) I hope you understood that!

OpenStudy (agentnao):

It might be a bit difficult for you to read typed up like this, but I can show you in a whiteboard room if you'd like.

OpenStudy (moongazer):

use first synthetic division if you know that

OpenStudy (anonymous):

@agentNao thank you so much ill take it from here :)

OpenStudy (agentnao):

You're very welcome! I hope I was of some help :)

OpenStudy (anonymous):

x = -1 is also a solution.

OpenStudy (anonymous):

Cubics have 3 solutions :)

OpenStudy (agentnao):

Oops! Sorry about that. Dangolberry is very right. When you square root something, you'll always have a plus and minus of the answer.

OpenStudy (anonymous):

so just to clarify, x=+-1 and +-2?

OpenStudy (anonymous):

x = 1, x = -1 and x = 2

OpenStudy (agentnao):

It's not +- 2 because we didn't have to square root that.

OpenStudy (anonymous):

right!